If a buffer solution is 0.150 M in a weak base (Kb = 1.0 × 10-5) and 0.430 M in its conjugate acid, what is the pH?
Answer #1

Kb = 1*10^-5

pKb = - log (Kb)

= - log(1*10^-5)

= 5

use:

pOH = pKb + log {[conjugate acid]/[base]}

= 5+ log {0.43/0.15}

= 5.457

use:

PH = 14 - pOH

= 14 - 5.4574

= 8.5426

Answer: 8.54